# MM-PhyRLHF: Reinforcement Learning Framework for Multimodal Physics Question-Answering

### v1

@@ Abstract.
Recent advancements in LLMs have shown their significant potential in tasks like text summarization and generation. Yet, they often encounter difficulty while solving complex physics problems that require arithmetic calculation and a good understanding of concepts. Moreover, many physics problems include images that contain important details required to understand the problem’s context. We propose an LMM-based chatbot to answer multimodal physics MCQs. For domain adaptation, we utilize the MM-PhyQA dataset comprising Indian high school-level multimodal physics problems. To improve the LMM’s performance, we experiment with two techniques, RLHF (Reinforcement Learning from Human Feedback) and Image Captioning. In image captioning, we add a detailed explanation of the diagram in each image, minimizing hallucinations and image processing errors. We further explore the integration of Reinforcement Learning from Human Feedback (RLHF) methodology inspired by the ranking approach in RRHF to enhance the human-like problem-solving abilities of the models. The RLHF approach incorporates human feedback into the learning process of LLMs, improving the model’s problem-solving skills, truthfulness, and reasoning capabilities, minimizing the hallucinations in the answers, and improving the quality instead of using vanilla-supervised fine-tuned models. We employ the LLaVA open-source model to answer multimodal physics MCQs and compare the performance with and without using RLHF.

### v2

@@ Abstract.
Recent advancements in LLMs have shown their significant potential in tasks like text summarization and generation. Yet, they often encounter difficulty while solving complex physics problems that require arithmetic calculation and a good understanding of concepts. Moreover, many physics problems include images that contain important details required to understand the problem’s context. We propose an LMM-based chatbot to answer multimodal physics MCQs. For domain adaptation, we utilize the MM-PhyQA dataset comprising Indian high school-level multimodal physics problems. To improve the LMM’s performance, we experiment with two techniques, RLHF (Reinforcement Learning from Human Feedback) and Image Captioning. In image captioning, we add a detailed explanation of the diagram in each image, minimizing hallucinations and image processing errors. We further explore the integration of Reinforcement Learning from Human Feedback (RLHF) methodology inspired by the ranking approach in RRHF to enhance the human-like problem-solving abilities of the models. The RLHF approach incorporates human feedback into the learning process of LLMs, improving the model’s problem-solving skills, truthfulness, and reasoning capabilities, minimizing the hallucinations in the answers, and improving the quality instead of using vanilla-supervised fine-tuned models. We employ the LLaVA open-source model to answer multimodal physics MCQs and compare the performance with and without using RLHF. **footnotetext: These authors contributed equally to this work
@@ 7. Conclusion
This study investigates the impact of incorporating image captioning models on the multimodal reasoning capabilities of open-source multimodal large language models (LLMs). Our findings indicate that providing the LLM with only the question text and the image (Table 1) results in the poorest performance. In contrast, augmenting the input with an image caption generated by a captioning model, along with the question text and image (Table 2), significantly improves reasoning accuracy. Interestingly, we observe that excluding the image entirely and supplying only the caption and question text (Table 3) yields worse performance than the comprehensive setting in Table 2, where the model is provided with the question text, image, and caption.
Future work will extend this approach by leveraging Reinforcement Learning with Human Feedback (RLHF). By incorporating human preference knowledge, we aim to align the reasoning capabilities of multimodal LLMs more closely with human expectations, further enhancing their performance and applicability in real-world scenarios.